for each of the cases below, show the appropriate t-value and the approximate normal z value. when a- confidence level=80%,df=40
and b- confidence level=99%,df=9
Answer #1

a. C = 0.80, df = 40

a-1-c= 1-0.80-0.20

t value :

teta/2dfo. 20/2,40 -1.303

( We look into t table across df = 40 and probability 0.20 for two tailed or 0.10 for one tailed and find t value across it )

Z value :

Z_{c}=t_{1-alpha /2}=t_{1-0.20/2}=1.28

( We look in to the probability section of normal probability table for the probability 0.90 and find z value across it )

b.

C = 0.99, df = 9

a-1-c= 1-0.99-0.01

t value :

c/2,dfo.01/2,9 3.250

( We look into t table across df = 9 and probability 0.01 for two tailed or 0.005 for one tailed and find t value across it )

Z value :

= ti-a/2 t1-0.01 /2 2.58

( We look in to the probability section of normal probability table for the probability 0.995 and find z value across it )